# Safe Explicable Policy Search

### v1

@@ IV Problem Formulation
In this section, we motivate and discuss the formulation of SEPS by incorporating the benefits of CMDP solution techniques (specifically, that of CPO [6]) with EPS [19]. Extending CPO to generate safe explicable policies is not straightforward due to the consideration of the objective over two different MDP models. We refer the reader to SEP [16], which discusses the challenges of solving a constrained optimization problem defined over two separate transition models. Similarly, extending EPS to generate safe explicable policies requires the specification of a reconciliation parameter, and a weighting parameter for a couple of entropy terms to scale the different measures with the reward appropriately. We propose a formulation for SEPS that systematically combines the capabilities of CPO and EPS while eliminating their limitations towards finding safe explicable policies.
Exhibiting explicable behavior is important for successful teaming. Examples of RL approaches that consider human feedback are [17, 18, 28, 3]. A popular objective for finding explicable plan or policy is to optimize an objective that linearly combines the agent’s rewards with an explicability measure [2, 3, 16]. However, being explicable at the cost of safety can be dangerous. Hence, it is unacceptable to compromise safety.
Ensuring safety in a task means avoiding or reducing harm. Implicitly, any task assumes successful completion without causing harm. For instance, if an agent is tasked with fetching hot coffee, an underlying safety assumption is that it does not bump into anyone on its way or spill the hot coffee. One approach, albeit nonideal, is to optimize an objective that balances risk penalties along with rewards, and explicability. This requires carefully scaling each component to provide a meaningful learning signal at every step, preventing issues like reward hacking. In safety-critical tasks, designing a single reward that simultaneously drives task completion and discourages unsafe behavior is challenging. A common approach is to define the reward as $\mathcal{R}=R-\omega C$, where $R$ represents the task reward component, and $C$ denotes the safety cost. However, results from CPO [6] demonstrate high sensitivity to the penalty coefficient $\omega$, highlighting the difficulty of crafting an effective overloaded reward function. Prior research [29], also emphasizes the importance of separating the task objective from safety constraints, as done in CMDPs, to enable the agent to learn meaningful behavior.
CMDP methods such as [5, 6, 7, 8], etc. efficiently optimize the task objective while enforcing safety constraints. Ideally, an agent should act both explicably (which involves considering two models) and efficiently while satisfying safety requirements. However, a CMDP requires both the objective and constraints to be defined under a single transition model. To address this, we propose to consider the objective in EPS, since it converts the consideration of two models to the consideration of two linearly combined objectives under a single transition model, as shown below,
@@
This objective has two main limitations, 1) similar to the penalty coefficient, determining an appropriate reconciliation factor ($\lambda$) can be challenging, and 2) since the task objective is maximized alongside explicability, for $\lambda>0$, there is no guarantee of task performance in the agent’s model, especially when $\lambda$ is high. To overcome these issues, we propose to constrain the agent’s task objective instead of linearly combining it with the explicability measure as below.
$\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{d}}$ $\mathcal{E}(\pi,\mathcal{M}_{A},\pi_{A}^{H},\mathcal{M}_{A}^{H})$
$s.t.\hskip 3.0pt$ $J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)\leq d_{i}\hskip 10.0pti=1,....,k$ (4)
where $\mathcal{E}(.)$ can be written as $J_{\mathcal{E},\mathcal{T}_{A}}(\pi)$ based on Eq. 2. Essentially, $J_{\mathcal{E},\mathcal{T}_{A}}(\pi)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}(u_{H}(s_{t},a_{t})+\mathcal{H}_{\mathcal{T}_{A}}(s_{t+1}|s_{t},a_{t})+\mathcal{H}_{\pi_{\theta}}(a_{t}|s_{t}))]$. Since the entropy terms ($\mathcal{H}_{\mathcal{T}_{A}}$ and $\mathcal{H}_{\pi_{\theta}}$) are approximated separately from the surrogate reward ($u_{H}$), they require additional scaling parameters which are not straightforward to determine. In our work, we mitigate this by simply maximizing $u_{H}$. This means that instead of maximizing the negative KL divergence between the trajectory distributions, we maximize the probability of the most expected trajectory that the agent can generate.
First, let us define the components involved in formulating SEPS. The agent’s decision model is modeled as a CMDP augmented with an additional limit on the task objective i.e., $\mathcal{M}_{A}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A},\mathcal{R}_{A},\{\mathcal{C}_{i}\}_{i=1:k},\{d_{i}\}_{i=0:k},\rho,\gamma\rangle$ where, $d_{0}$ is the limit on the task objective $\mathcal{R}_{A}$ and $d_{1},....,d_{k}$ are limits on the cost functions $\mathcal{C}_{1},....,\mathcal{C}_{k}$, respectively. The user’s belief of the agent’s model is modeled as an MDP i.e., $\mathcal{M}_{A}^{H}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A}^{H},\mathcal{R}_{A}^{H},\rho,\gamma\rangle$. Given the two decision models, we define SEPS as below.
@@
Safe Explicable Policy Search (SEPS) is the problem of searching for a policy that maximizes the probability of the user’s expected trajectory subject to the constraints on the task objective and the costs associated with safety in the agent’s model, or formally,
@@
$s.t.\hskip 3.0ptJ_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)$ $\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)$ $\leq d_{i}\hskip 10.0pti=1,...,k.$ (5)
It is interesting to note that in the special case where $\mathcal{T}_{A}$ and $\mathcal{T}_{A}^{H}$ are identical, $u_{H}$ closely approximates the user’s reward model $\mathcal{R}_{A}^{H}$. However, we note that the goal of learning $u_{H}$ is to obtain feedback on the user’s assessment of explicableness (based on his own model - $\mathcal{T}_{A}^{H}$, $\mathcal{R}_{A}^{H}$) of agent’s behavior, i.e., when the agent operates under the true domain dynamics $\mathcal{T}_{A}$. Therefore, the SEPS problem in Eq. 5 is equivalent to the Safe Explicable Planning problem [4]
@@
$s.t.\hskip 3.0ptJ_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)$ $\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)$ $\leq d_{i}\hskip 10.0pti=1,...,k.$ (6)
V Safe Explicable Policy Search
@@
A previous study on Safe Explicable Planning (SEP) [4] proposes searching through the policy space while evaluating policies in the agent’s model for feasibility and in the user’s model for explicablity. The SEP algorithms iterate over all states and actions to find an optimal, safe explicable policy, which is impractical in continuous MDPs. The SEPS problem in Eq. 5 reduces to a CMDP, allowing us to use its policy search techniques. CPO [6] is a popular approach to solving CMDP as it ensures safe behavior even during the learning process. However, the algorithm is limited to solving an objective with a single constraint. Therefore, we require an analytic solution to efficiently solve SEPS. As the number of constraints increases, deriving feasible bounds for the solution becomes challenging due to the intersection of feasible regions of every constraint.
Let us introduce some useful notations. The on-policy value function and action-value function in the agent’s model, over a general reward function $\mathcal{R}$, are given by $V_{\mathcal{R}}^{\pi}(s)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}{=}s]$ and $Q_{\mathcal{R}}^{\pi}(s,a)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}{=}s,a_{0}{=}a]$, respectively. Then, the corresponding advantage function is $A_{\mathcal{R}}^{\pi}(s,a)\dot{=}Q_{\mathcal{R}}^{\pi}(s,a)-V_{\mathcal{R}}^{\pi}(s)$. The function $\mathcal{R}$ may be replaced by $u_{H}$, $\mathcal{R}_{A}$, or any $\mathcal{C}_{i}$.
CPO applies trust region policy optimization (TRPO) [29] to policy search updates in a CMDP. Using the notation above, the formulation in Eq. 5 can be written as a policy update rule given by
[…]
where, $\Pi_{\theta}\subset\Pi$ is a set of parameterized policies with parameters $\theta$, $\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})=\mathbb{E}_{s\sim\pi^{n}}[\mathcal{D}_{KL}(\pi||\pi_{n})[s]]$ is a measure of KL Divergence between the current policy $\pi_{n}$ and a potential new policy $\pi$, and $\delta>0$ is the step size. The set of policies that satisfy the third constraint is called the trust region, i.e. $\{\pi_{\theta}\in\Pi_{\theta}|\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})\leq\delta\}$. For performance guarantees and constraint satisfaction guarantees, see [29] and [6], respectively.
@@
For continuous MDPs with high dimensional spaces, solving Eq. 7 exactly is intractable and requires an approximation. The CPO method proposes the following approximation
$\theta_{n+1}=\operatorname{arg\,max}_{\theta}g^{T}(\theta-\theta_{n})\\ s.t.\hskip 3.0ptc_{0}-b_{0}^{T}(\theta-\theta_{n})\leq 0\hskip 70.0pt\\ c_{i}+b_{i}^{T}(\theta-\theta_{n})\leq 0\hskip 10.0pti=1,...,k\\ \frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta\hskip 80.0pt$ (8)
where, $g$ is the gradient of the objective (explicability measure), $b_{0}$ is the gradient of the constraint on agent’s task objective, $b_{i}$ is the gradient of the constraint $i$, $c_{0}\dot{=}-J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi_{n})+d_{0}$, $c_{i}\dot{=}J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi_{n})-d_{i}$, and $H$ is the Hessian of the KL Divergence. The optimization problem mentioned above is convex. For policies with high-dimensional parameter spaces like neural networks, the primal problem above is impractical to solve (even with few constraints) using any convex program solver due to the computation complexity of the Hessian matrix.
Feasible case: When the problem is feasible the dual to the above is easier to solve and is given by
$\max_{\lambda\geq 0,\nu\succeq 0}\frac{-1}{2\lambda}(g^{T}H^{-1}g+2r^{T}\nu+\nu^{T}S\nu)+\nu^{T}c-\lambda\delta$ (9)
where $c\dot{=}[c_{0},c_{1},...,c_{k}]^{T}$, $B\dot{=}[b_{0},b_{1},...,b_{k}]$, $r\dot{=}g^{T}H^{-1}B$, and $S\dot{=}B^{T}H^{-1}B$. Here, the product terms involving the inverse of the Hessian matrix can be approximately computed as in [29]. If $\lambda^{*}$ and $\nu^{*}$ are optimal solutions to the dual above. The update rule for the feasible case is
$\theta_{n+1}=\theta_{n}-\frac{1}{\lambda}H^{-1}\left(g+\nu B\right)$ (10)
Infeasible case: When the policy $\pi_{\theta_{n}}$ is infeasible, due to a bad step taken previously (occurs due to approximation), the system can recover by updating the policy parameters to strictly reduce the constraint violation. In this case, the update rule is found by solving the following optimization problem depending upon the constraint being violated
$\theta_{n+1}=\operatorname{arg\,min}_{\theta}\hskip 5.0ptc_{m}-b_{m}^{T}(\theta-\theta_{n})\hskip 15.0pt\text{where, }m\in[0,k]\\ s.t.\,\,\,\,\frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta\hskip 50.0pt$ (11)
The dual to the above is
$\max_{\phi\geq 0}-\frac{s}{2\phi}+c_{m}-\phi\delta.$ (12)
where $s=b_{m}^{T}H^{-1}b_{m}$. If $\phi^{*}$ is the solution to the dual above. The update rule for the infeasible case is
$\theta_{n+1}=\theta_{n}-\frac{1}{\phi}H^{-1}b_{m}$ (13)
When multiple constraints are violated, we solve Eq. 11 for each violating constraint and combine the solutions in Eq. 13. This adjusts the policy towards the feasible region w.r.t. all constraints in a single step. Alternatively, one could update the policy with respect to only one constraint at a time.
We derive an analytical solution to SEPS for both feasible and infeasible cases above (refer to Theorems 1 and 2 in the extended version of this paper) when a single cost function is present, resulting in two constraints.
VI Evaluation
We evaluated our method in continuous safety-gym environments [30] and with a physical robot experiment. We designed the evaluation scenarios, including the risk elements. In all of the evaluations, we assume a learned surrogate reward function that captures SEPS relevant information regarding the user model. The goal of our evaluation is 1) to validate the efficacy of our method to search for safe explicable policies which a) meet user expectations, and b) are safe and efficient, 2) to analyze the behaviors generated by our method in comparison with that in the agent’s model and user expectations, and 3) to demonstrate the applicability of SEPS to real-world scenarios.
VI-A Safety
In our first evaluation, we aim to answer ‘What are the benefits of safety constraints in explicable planning?’ and ‘Is SEPS’s policy explicable under safety constraints?’. We design an evaluation scenario such that the user’s expectations conflict with the safety requirements of the agent.

### v3

@@ IV Problem Formulation
To formulate safe explicable policy search (SEPS), similar to EPS, the user’s belief of the agent’s model is modeled as an MDP i.e., $\mathcal{M}_{A}^{H}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A}^{H},\mathcal{R}_{A}^{H},\rho,\gamma\rangle$; furthermore, the agent’s decision model is modeled as a CMDP i.e., $\mathcal{M}_{A}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A},\mathcal{R}_{A},\{\mathcal{C}_{i}\}_{i=1:k},\{d_{i}\}_{i=0:k},\rho,\gamma\rangle$. To focus on policy search, we assume access to the surrogate reward function $u_{H}$ in EPS.
IV-A Rationale
Explicability and safety are orthogonal issues: an explicable policy may place too much focus on user understandability, resulting in unsafe behaviors when the user is simply unaware of the safety risks; on the other hand, an explicable policy may align well with safety if the user understands the risks but is risk-averse, resulting in overly conservative behaviors that negatively impact task efficiency. Hence, it is imperative for explicable behavior generation to be guarded by both safety constraints and efficiency criteria. Hence, it is imperative for explicable behavior generation to be guarded by safety constraints. However, we will explain next that a few alternatives would not work well.
IV-A1 Multi-objective optimization with linear combination
One idea is to optimize an objective that balances safety costs with rewards and explicability. The corresponding optimization problem can then be defined as maximizing $\mathcal{R}^{\prime}=\mathcal{R}-\omega_{1}\mathcal{C}+\omega_{2}\mathcal{E}$, where $\omega_{1}$ and $\omega_{2}$ are the respective scaling parameters.
IV-A2 Constrained EPS
Alternatively, we can consider EPS subject to safety constraints as follows:
@@ IV-A2 Constrained EPS
The above two alternative formulations considered share two related limitations: 1) determining an appropriate reconciliation/weight factors can be difficult. Designing a single reward that simultaneously drives task completion, meets user expectations, and discourages unsafe behavior is challenging. It requires carefully scaling each component to provide a meaningful learning signal at each step, preventing issues like reward hacking. For example, in CPO [6], the results of optimizing $\mathcal{R}-\omega_{1}\mathcal{C}$ demonstrate a high sensitivity to the cost coefficient $\omega_{1}$, highlighting the difficulty of crafting an effective overloaded reward function. Prior research [29], also emphasizes the importance of separating the task objective from safety constraints, as in CMDPs [5, 6, 7, 31], to enable the agent to learn meaningful behavior. 2) since the task objective is maximized in combination with explicability, for $\lambda>0$, there is no guarantee of task performance or even achievement, since the agent may be able to increase the explicability score to “compensate” for the reduction in task performance.
IV-B Safe Explicable Policy Search (SEPS)
In SEPS, we observe that the two components of EPS can be separated to remove the reconciliation factor and let the agent reward be absorbed into the safe constraints in CMDP, with the limit $d_{0}$. Even though this would introduce a constraint parameter, it has an intuitive appeal: the suboptimality level of the behavior, which can be used to ensure the task performance/achievement as desired. Hence, such a parameter can be set more straightforwardly by the engineer than the semantically confusing reconciliation factor.
A remaining challenge is the additional entropy terms in $\mathcal{E}$. Since the surrogate reward $u_{H}$ in EPS is learned to reflect $\mathcal{H}_{\mathcal{T}_{A}^{H}}$ and $\mathcal{H}_{\pi_{A}^{H}}$, combining it with the remaining entropy terms ($\mathcal{H}_{\mathcal{T}_{A}}$ and $\mathcal{H}_{\pi_{\theta}}$) requires additional scaling parameters. In SEPS, we simplify by removing these terms. This means that instead of maximizing the negative KL divergence between the trajectory distributions, we maximize the probability of the most expected trajectory that the agent can generate. This leads to an approximation of the explicability objective that encourages a narrower distribution than the original objective.
@@ IV-B Safe Explicable Policy Search (SEPS)
Safe Explicable Policy Search (SEPS) is the problem of searching for a policy that maximizes the probability of the user’s expected trajectory, captured by $u_{H}$, subject to the constraints on the task objective $\mathcal{R}_{A}$ and the costs $\mathcal{C}_{i}$ associated with safety in the agent’s model, or formally,
@@
$s.t.\hskip 3.0pt$ $\color[rgb]{0,0,0}{C_{0}:\,\,\,\,}J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)\geq d_{0}$
$\color[rgb]{0,0,0}{C_{i}:\,\,\,\,}J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)\leq d_{i}\hskip 10.0pti=1,...,k.$ (6)
The main objective here is to maximize $u_{H}$ under the true domain dynamics $\mathcal{T}_{A}$, which in turn maximizes the approximate explicability score. Although the SEPS problem in Eq. (6) appears similar to the SEP [4] problem below
@@
$s.t.\hskip 3.0ptJ_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)$ $\geq d_{0},$ (7)
there are a few critical differences: 1) in SEPS, the constraint on the agent’s reward or task performance, whereas it was associated with the safety constraint in [4]. It requires the safety constraint to be correlated with task performance, which is not always feasible; 2) in [4], $\mathcal{M}_{A}^{H}$ is assumed to be given whereas it was replaced here by the surrogate reward function that can be learned, which is evaluated under the true domain dynamics. Such a subtle but important difference allows us to apply learning methods to solve the problem instead of replying on planning; 3) formulating SEPS as a CMDP allows us to leverage CMDP solutions to handle multiple constraints in continuous domains, instead of as a complex search problem under two different models in [4].
V Search Method
@@
When the MDP models are known, as in Eq. (7), SEP [4] proposes searching through the policy space while evaluating policies in the agent’s model for feasibility and in the user’s model for explicablity. The SEP algorithms iterate over all unique deterministic policies (formed by iterating over all states and actions) to find an optimal safe, explicable policy, which is impractical in continuous MDPs. The SEPS problem in Eq. (6) reduces to a CMDP, allowing us to use its policy search techniques.
We choose the CPO [6] algorithm to solve SEPS as it ensures safe behavior even during the learning process. However, the CPO algorithm is limited to solving an objective with a single constraint and does not provide the analytic solution for more than one constraint. As the number of constraints increases, deriving feasible bounds for the solution becomes challenging because of the intersection of feasible regions of every constraint. In this work, we adopt the CPO solution technique, but derive a new analytical solution for SEPS that includes at least two constraints.
Applying the CPO update rule to SEPS in Eq. (6) will yield an expression similar to that in Eq. (1). An approximation to that update rule (Eq. (2) as in CPO) is given by the primal below:
$\theta_{n+1}=\operatorname{arg\,max}_{\theta}g^{T}$ $(\theta-\theta_{n})$
$s.t.\hskip 3.0ptc_{0}-b_{0}^{T}(\theta-\theta_{n})$ $\leq 0$
$c_{1}+b_{1}^{T}(\theta-\theta_{n})$ $\leq 0$
$\frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})$ $\leq\delta$ (8)
where, $b_{0}$ is the gradient of the constraint on agent’s task objective, and $c_{0}\dot{=}-J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi_{n})+d_{0}$. $b_{1}$ and $c_{1}$ are as defined in Eq. (2).
The optimization problem mentioned above is convex. It is worth noting that for policies with high-dimensional parameter spaces like neural networks, the primal problem (even with few constraints) is impractical to solve using a convex program solver due to the computation complexity of the Hessian matrix, which motivates the need for an efficient analytical solution.
Feasible case: We consider the feasible case when the current policy $\pi_{\theta_{n}}$ satisfies all constraints. In this case, it is easier to solve the dual of Eq. (8), given by
$\max_{\begin{subarray}{c}\lambda\geq 0\\ \nu_{0}\geq 0\\ \nu_{1}\geq 0\end{subarray}}g^{T}x+\lambda(\frac{1}{2}x^{T}Hx-\delta)+\nu_{0}(b_{0}^{T}x+c_{0})+\nu_{1}(b_{1}^{T}x+c_{1})$ (9)
where $\lambda$, $\nu_{0}$, and $\nu_{1}$ are Lagrangian variables. The solution to the feasible case is given by Theorem 1 in Appendix A of the extended version of this paper [41]. To determine the optimal values of the Lagrangian variables, we need to consider the intersection of the trust region with the two linear constraints. Specifically, the optimal value of $\lambda$ depends on the following four cases, viz., 1) both linear constraints are active ($\nu_{0}>0$ and $\nu_{1}>0$), 2) only the first linear constraint is active ($\nu_{0}>0$ and $\nu_{1}=0$): which means that the trust region lies well within the feasible region of the second constraint and can be safely ignored, 3) only the second linear constraint is active ($\nu_{0}=0$ and $\nu_{1}>0$), and 4) both linear constraints are inactive ($\nu_{0}=0$ and $\nu_{1}=0$): which means that the optimization depends solely on the objective ($g^{T}x$). In brief, if $\lambda^{*}$, $\nu_{0}^{*}$, and $\nu_{1}^{*}$ are optimal solutions to the dual above, then, the update rule for the feasible case is
$\theta_{n+1}^{*}=\theta_{n}+\frac{1}{\lambda^{*}}H^{-1}\left(g+\nu_{0}^{*}b_{0}-\nu_{1}^{*}b_{1}\right)$ (10)
Infeasible case: We consider the infeasible case when the current policy $\pi_{\theta_{n}}$ violates one or both linear constraints. This could be the result of a bad step taken previously (occurs due to errors in approximation). The system can recover by updating the policy parameters to strictly reduce the constraint violation. In other words, the system must focus on satisfying the constraint(s) and ignore optimizing the objective ($g^{T}x$), temporarily. In this case, depending on the linear constraint that is violated in Eq. (8), the update rule is found by solving the primal below
$\theta_{n+1}=\operatorname{arg\,min}_{\theta}\hskip 5.0ptc_{m}+b_{m}^{T}(\theta-\theta_{n})\hskip 15.0pt\text{where, }m\in\{0,1\}\\ s.t.\,\,\,\,\frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta,\hskip 50.0pt$ (11)
and its corresponding dual is given by
$\max_{\phi\geq 0}-\frac{b_{m}^{T}H^{-1}b_{m}}{2\phi}+c_{m}-\phi\delta.$ (12)
where $\phi$ is a Lagrangian variable. The solution to the feasible case is given by Theorem 2 in Appendix A of the extended version of this paper [41].
In brief, if $\phi^{*}$ is the solution to the dual, then the update rule for the infeasible case is
$\theta_{n+1}^{*}=\theta_{n}-\frac{1}{\phi^{*}}H^{-1}b_{m}$ (13)
When multiple constraints are violated, we solve Eq. (11) for each constraint that is violated and linearly combine the solutions in Eq. (13). This adjusts the policy towards the feasible region w.r.t. all constraints in a single step. Alternatively, one could update the policy with respect to a single constraint at a time.
VI Evaluation
We evaluated our method in continuous Safety-Gym environments [42] and through a physical robot experiment, designing scenarios that included explicit risk elements. Across all evaluations, we assumed a learned surrogate reward function that captures sufficient information about user expectations in $\mathcal{M}_{A}^{H}$. Our goal is to 1) validate the ability of our method to find safe, efficient, and explicable policies that meet user expectations, 2) analyze the behaviors generated by our method relative to state-of-the-art and baseline approaches, and 3) demonstrate the applicability of SEPS in real-world scenarios.
VI-A Safety
In our first evaluation, our goal is to answer ‘What are the benefits of safety constraints in explicable planning?’ and ‘Can SEPS identify an explicable policy under safety constraints?’. We use the Safety-Gym PointGoal task (shown in Fig. 3(a)) where we assume user expectations conflict with the agent’s safety requirements. In $\mathcal{M}_{A}$, the agent must reach the goal (green) without entering hazardous regions (dark blue). The reward $\mathcal{R}_{A}$ is positive for progress toward the goal and negative for moving away, and $\mathcal{C}_{1}$ penalizes entry into hazardous regions. In contrast, according to user expectations ($u_{H}$), the agent can traverse hazardous regions as the user is unaware of their risks but must avoid the light blue boxes as the user regards them fragile.
@@ VI-A Safety
We evaluated SEPS against three baselines: AGT (optimal agent behavior), HUM (human expected behavior), and EPS. AGT optimizes $\mathcal{R}_{A}$, HUM optimizes $u_{H}$, and EPS optimizes the linear combination $\mathcal{R}_{A}+\lambda u_{H}$ (see the Preliminary Section). All baselines were trained with PPO [43]. Here, the need for safety constraints in explicable planning is highlighted, as the baselines do not incorporate them. We compared the resulting behaviors and measured the returns under the surrogate reward, the agent’s task reward, and the cost function.
(a) PointGoal
(b) PointButton
[…]
Results and Discussion. Figure 3(a) illustrates the behaviors generated. AGT reached the goal via the shortest path by moving one of the boxes, contrary to user expectations. HUM and EPS avoided the boxes, but entered hazardous regions due to a lack of safety constraints. In contrast, SEPS efficiently reached the goal while avoiding both boxes and hazardous regions.
These behaviors aligned with the quantitative results in Fig. 2. The primary finding is that all baselines violated safety constraints, as reflected by returns in $\mathcal{C}_{1}$ (Fig. 2(c)), where values farther from zero indicate greater violations. AGT, HUM, and EPS showed higher violations as their objectives ignore safety specifications. AGT also performed poorly on user expectations (Fig. 2(a)), prioritizing the shortest path even if it involves moving boxes. HUM achieved high returns in $\mathcal{R}_{A}$ (Fig. 2(b)) as it does not deviate much from the goal. EPS, despite using the best reconciliation factor ($\lambda=2$), showed instability in learning $\mathcal{R}_{A}$ (Fig. 2(b)) due to its linear combination objective. SEPS outperformed the baselines on the agent’s task (Fig. 2(a)) and user expectations (Fig. 2(b)) while satisfying safety requirements (Fig. 2(c)).
VI-B Ablation Study
In our second evaluation, our goal is to answer ‘Can safe explicable policy search be formulated differently?’ or, equivalently, ‘What is the benefit of the SEPS formulation?’. We used the Safety-Gym PointButton task (Fig. 3(b)), where the user’s model aligns with the agent’s safety requirements but differs from its task objective.